OpenStudy (anonymous):

Integration Help ; How to integrate combination of exponential and trignometric . If I use part by part integration method using u and v , exponential integration will outputs the same result . Help me....

OpenStudy (anonymous):

It can be the same, it depends on which function you are given ...

OpenStudy (anonymous):

can you post an example

OpenStudy (anonymous):

Okay.....w8...

OpenStudy (anonymous):

just keep doing it by parts until you get the same expression on the RHS as the one on the LHS so you will have something like F= blabla +blala2 + blabla3 + something*F. now just solve for F

OpenStudy (anonymous):

where F is the original integral...

OpenStudy (anonymous):

Please see my attached image....

OpenStudy (anonymous):

is that e^-x or e^-n?

OpenStudy (anonymous):

Its e^-x . Sorry for the bad handwriting .

OpenStudy (anonymous):

let F= \[\int\limits_{0}^{1}e^{-x}\cos(n \pi x)dx\]. Now try doing integration by parts with u and v (for every integral use the same substitution for u and v ). You will end up with "something" + the original integral after few iterations. And if you just write F insted of your original integral you will have 2F=some_expression_here1 + some_expression_here2 + some_expression_here3 + some_constant*F

OpenStudy (anonymous):

now just solve for F, since F = original integral

OpenStudy (anonymous):

I will try to do a sketch and show you . Please check it .

OpenStudy (anonymous):

I've done partial solution . I have done until 2 times part by part . Its just looking like gonna continue forever . Please check and advice .

OpenStudy (anonymous):

ok, if you take out 1/pi*n in the very last integral you will get the original integral, that is F. Now, that whole row (last 1) is equal to F...now you have an algebraic equation, (solve for F)

OpenStudy (anonymous):

i didn't check for any errors, i just checked the method...:D

OpenStudy (anonymous):

Here's my solution hopefully there's no errors. I had a quick look. Good luck.

OpenStudy (anonymous):

In the fifth line, there shouldn't be a 2 infront of the n^2 pi^2 for my solution, the rest is good though.

OpenStudy (anonymous):

Please check my image .

OpenStudy (anonymous):

Okay now all you have to do is plug in the values \[2 \int \limits_{0}^{1} e^{-x}cos{n\pi x} dx =\frac{ 2e^{-x}(n\pi sin{n\pi x} - cos{n\pi x})}{n^2\pi^2 +1}\]\[=\frac{ 2e^{-1}(n\pi sin{n\pi} - cos{n\pi })}{n^2\pi^2 +1}+\frac{ 2}{n^2\pi^2 +1}\]\[=\frac{ 2(n\pi sin{n\pi} - cos{n\pi })}{en^2\pi^2 +e}+\frac{ 2e}{en^2\pi^2 +e}\]\[=\frac{ 2(n\pi sin{n\pi} - cos{n\pi }+e)}{en^2\pi^2 +e}\]

OpenStudy (anonymous):

okay yes I did bring that integral to the other side and added it, so let's look at just the numbers infront of the integral. you will have \[1+\frac{1}{n^2\pi^2}=\frac{n^2\pi^2+1}{n^2\pi^2}\]

OpenStudy (anonymous):

How you got 1 + 1/n^2.pi^2 . I got no idea how u got the 1 .

OpenStudy (anonymous):

the one comes from the original integral. \[\int e^{-x}cos{n\pi x} = ....... -\frac{1}{n^2\pi^2}\int e^{-x}cos{n\pi x}\] \[1\int e^{-x}cos{n\pi x} +\frac{1}{n^2\pi^2}\int e^{-x}cos{n\pi x}= ....... \] \[(1+\frac{1}{n^2\pi^2})\int e^{-x}cos{n\pi x}= ....... \] \[(\frac{n^2\pi^2+1}{n^2\pi^2})\int e^{-x}cos{n\pi x}= ....... \]

OpenStudy (anonymous):

Okay.....now i got u....thanks